1. In the chemical reaction below, determine how much oxygen is required to burn 2.8 g of propane with complete combustion.



C3H8 + 5 O2 à 3 CO2 + 4 H2O







2. 2 g of Magnesium is placed into a solution of Copper(I) Chloride> Write the proper balanced chemical reaction with states. Predict the mass of Copper that could be produced.
1
Expert's answer
2020-04-29T13:29:29-0400
  1. Proportion: m(C3H8)/Mr(C3H8)=m(O2)/5*Mr(O2)

m(O2)=m(C3H8)*5*Mr(O2)/Mr(C3H8)=2.8*5*32/44=10.18 g

2.Mg(s) + CuCl2 (aq)--> Cu(s) + MgCl2(aq)

m(Cu)=m(Mg)*Mr(Cu)/Mr(Mg)=2*64/24=5.33 g